Opened 6 years ago

Last modified 6 years ago

#421 assigned flight processing

BGS08/03, flight day 271/2011, Glasgow

Reported by: knpa Owned by: knpa
Priority: alpha 4 high Milestone: 2011 data processing completion
Component: Archiving Keywords:
Cc: Other processors: jaho,knpa

Description (last modified by adbe)

Data location: ~arsf/arsf_data/2011/flight_data/uk/BGS08_03-2011_271_Glasgow

Data arrived from ARSF via network transfer on 06/10/2011

BGS08/03 Scientific objective:

  • To identify sediment character of tidal flats
  • To identify contaminated and industrial waste

BGS08/03 Priority: a4h

PI: Colm Jordan

Notes:
Collection had to be stopped early due to deteriorating weather.

Sensors:

Camera (07/11/2011)
Eagle (07/11/2011)
Hawk (07/11/2011)
LiDAR (07/11/2011)

Change History (22)

comment:1 Changed 6 years ago by knpa

  • Description modified (diff)

comment:2 Changed 6 years ago by adbe

Started to process navigation

comment:3 Changed 6 years ago by adbe

Processed Navigation

comment:4 Changed 6 years ago by jaho

Started eagle and hawk processing.

comment:5 Changed 6 years ago by adbe

Started processing LiDAR

comment:6 Changed 6 years ago by knpa

  • Priority changed from alpha 4 medium to alpha 4 high

Starting RCD processing.

comment:7 Changed 6 years ago by jaho

Putting SCT correction on hold until LiDAR is processed so that it can be verified against an intensity image.

comment:8 Changed 6 years ago by adbe

Corrected LiDAR for pitch error

comment:9 Changed 6 years ago by adbe

LiDAR ready for delivery check

comment:10 Changed 6 years ago by jaho

Hyperspectral ready for delivery check.

comment:11 Changed 6 years ago by jaho

LiDAR delivery check complete.

Can't open files with LAG as it crashes with a segfault. Possibly due to files size?

Otherwise everything looks good.

comment:12 Changed 6 years ago by knpa

Gary says this PI should have access to their own vectors so do not wait for these before delivering.

comment:13 Changed 6 years ago by adbe

Hyperspectral delivery check:

  • The dem header says that it is in lat/lon, but the dem is called *_utm.dem, and if it was formed from lidar should be called BGS08_03-2011_271-lidar_ASTER.dem
  • The DEM name in the aplcorr command is BGS08 03-2011-271-lidar_NextMap.dem, should be BGS08_03-2011_271-ASTER.dem
  • Probably need to include misc/OSTN02_NTv2.gsb in the delivery 'used in apltran command', this file is in users/rsg/arsf/dems/ostn02/OSTN02_NTv2.gsb
    • You will probably need to change the ReadMe delivery contents
    • Change the apltran command
  • Need to change the quality issues to say just eagle first and last bands are blank

I'll probably need to do a check for the apl commands after these changes have been made.

comment:14 Changed 6 years ago by jaho

Changed the DEM, added misc/OSTN02_NTv2.gsb and updated the ReadMe.

comment:15 Changed 6 years ago by adbe

Hyperspectral apl commands work, ready to deliver

comment:16 Changed 6 years ago by knpa

RCD ready for delivery check.

comment:17 Changed 6 years ago by adbe

RCD Delivery Check:

*BGS08_03-Glasgow.kml should be BGS08_03-2011_271.kml

Otherwise ready to deliver.

comment:18 Changed 6 years ago by knpa

Fixed, ready to deliver.

adbe is preparing Glasgow disk to go out.

comment:19 Changed 6 years ago by adbe

  • Description modified (diff)

Delivered Camera, Hyperspectral and LiDAR to Colm Jordan on 07/11/2011 on disk 118

comment:20 Changed 6 years ago by mark1

Hyperspectral XML files have been re-created and re-delivered to PI (via email). This corrects Eagle IFOV value in XML and extra comments included.

comment:21 Changed 6 years ago by adbe

Hyperspectral sct's

lineeaglehawk
1-0.09-0.05
2-0.02-0.06

comment:22 Changed 6 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from adbe to knpa
  • Status changed from new to assigned

Beginning archiving.

Note: See TracTickets for help on using tickets.